Circular No. 7705

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ee IN NGC 2347 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong of an apparent supernova (mag 17) on an unfiltered CCD image (limiting mag 19) taken on Sept. 1.160 UT in the manner noted on IAUC 7702 (SN 2001eb). From this image, SN 2001ee was measured at α = 7^{h}16^{m}04^{s}.12, δ = +64^{o}42′19′′.7 (equinox 2000.0), which is 5′′ west and 20′′ south of the center of NGC 2347. Images taken by Armstrong on Sept. 5.015 and 5.127 show the new object at mag 17.1 and yield averaged position end figures 04^{s}.14, 19′′.6. Nothing is visible at this position on an image obtained by Armstrong on 2000 Oct. 2 (limiting mag 19.5), and nothing is present on second Palomar Sky Survey images from 1989 Dec. 23 (red, limiting mag 20.8), 1997 Mar. 9 (blue, limiting mag 22.5), and 1995 Nov. 16 (infrared). Hurst adds that SN 2001ee is absent from an image in J. C. Vickers’ Deep Space CCD Atlas: North (1993, p. 75; limiting mag ∼ 19.5).

SUPERNOVA 2001du IN NGC 1365 S. D. Van Dyk, Infrared Processing and Analysis Center, California

Institute of Technology; and W. D. Li and A. V. Filippenko, University of California at Berkeley, write: “We have attempted identification of the progenitor of SN 2001du (IAUC 7690, 7692, 7704) on Hubble Space Telescope (HST) archive-image data originally obtained by GO-5222 (unfortunately, the much deeper Key Project data obtained by GO-5972 do not contain the site of this supernova). We have remeasured an accurate position from an image provided by G. Bock, astrometrically reconstructed to ACT stars: α = 3^{h}33^{m}29^{s}.10, δ = −36^{o}08′31′′.2 (equinox 2000.0). From this position, allowing a conservative error of 1′′.5, we isolate three possible progenitor stars on the F814W image, with magnitudes I ≈ 23.1, 23.4, and 23.3 (± 0.5), in order of increasing distance from our position. Only the third star (with V ≈ 23.9 ± 0.5 and V–I = +0.59) actually has a detected counterpart on the F555W image. The star closest to the position has a lower limit to its V–I color of ∼ +0.85 mag. Assuming a distance modulus of 31.31 mag, this star has an absolute I magnitude of −8.2, which is not inconsistent with what would be expected for a late-type (red) supergiant star. If the progenitor was not one of these three stars, we can only place an upper limit on its absolute brightness at I ≈ −7.0. Clearly, images of the supernova at later times, when it has greatly faded, must be obtained, preferably with HST, to better pinpoint the supernova in its environment and to help constrain its nature based on its stellar and gaseous environment.”

2001 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green